# RSA: Resolving Scale Ambiguities in Monocular Depth Estimators through Language Descriptions

### v1

@@ Abstract
We propose a method for metric-scale monocular depth estimation. Inferring depth from a single image is an ill-posed problem due to the loss of scale from perspective projection during the image formation process. Any scale chosen is a bias, typically stemming from training on a dataset; hence, existing works have instead opted to use relative (normalized, inverse) depth. Our goal is to recover metric-scaled depth maps through a linear transformation. The crux of our method lies in the observation that certain objects (e.g., cars, trees, street signs) are typically found or associated with certain types of scenes (e.g., outdoor). We explore whether language descriptions can be used to transform relative depth predictions to those in metric scale. Our method, RSA, takes as input a text caption describing objects present in an image and outputs the parameters of a linear transformation which can be applied globally to a relative depth map to yield metric-scaled depth predictions. We demonstrate our method on recent general-purpose monocular depth models on indoors (NYUv2) and outdoors (KITTI). When trained on multiple datasets, RSA can serve as a general alignment module in zero-shot settings. Our method improves over common practices in aligning relative to metric depth and results in predictions that are comparable to an upper bound of fitting relative depth to ground truth via a linear transformation.
@@ 1 Introduction
We consider whether an additional modality can be used to resolve the scale ambiguity in single-image 3D reconstruction, i.e., transforming scaleless relative depth to metric depth. One might observe that natural (including man-made) scenes do not occur by chance, but rather by design, with the regularity of object-scene co-occurrences [18, 27]: Certain scenes (e.g., outdoors) are composed of certain categories of objects (e.g., cars, trees, buildings) and associated with a certain order of magnitude in scale (e.g. tens of meters). Hence, we hypothesize that language, in the form of text captions or descriptions, can be used to infer the scale of the 3D scene and to transform relative depth to metric depth. The choice of language also has practical value in that it does not require costly data acquisition with an additional synchronized and calibrated sensor (e.g., lidar, time-of-flight). With the availability of publicly available pre-trained panoptic segmentation and object detection models and image captioners, one can automate the data acquisition, training, and inference process. Nonetheless, they are not a necessary part of the work, but may facilitate ease of use.
@@
To test the feasibility of our hypothesis, we consider monocular depth estimation, where a strong prior is necessary for inference; this prior may come from an image, or an independent modality such as language. Specifically, we consider monocular depth models belonging to the general-purpose, relative depth estimation paradigm to control for side effects such as the scale learned along with shape in existing works that focus on predicting metric depth for a specific dataset. To this end, we assume access to a pre-trained scaleless monocular depth estimation model; aside from a set of images, we also assume text captions describing them. As the standard procedure to factor out scale is to normalize through a linear transformation, we propose to learn a parameterized function that predicts the parameters of a linear transformation based on the text description. Applying them to a relative depth map and inverting its values will transform it to a metric-scaled depth map to resolve the scale ambiguity. We term our method RSA, as an acronym for “Resolving Scale Ambiguities”.
@@
In one mode, like existing works that also transform relative to metric depth, albeit with images and domain-specific scaling [3], we train specific models of RSA for specific datasets (e.g., NYUv2 [45], KITTI [15]). In another mode, when trained on multiple datasets across different domains, RSA generalizes well and can not only handle images sampled from the datasets it was trained on, but also those from novel datasets in a zero-shot manner. The use of language, which is invariant to illumination, object orientation, occlusion, scene layout, etc., many of the nuisance variables that vision algorithms are sensitive to, demonstrates a promising avenue for general-purpose relative to metric scale recovery to complement the growing works in monocular depth estimation. We evaluate our method on indoor and outdoor benchmarks, where we improve over common practices in aligning relative depth to metric scale. We show that using RSA is comparable to matching relative depth via a linear transformation to the ground truth, which is the performance upper bound.
@@ 2 Related Work
Monocular depth estimation using relative depth. Trained across different domains, self-supervised depth estimators trained with multi-view photometric objects produce up-to-scale predictions that are linearly correlated with their absolute depth values across the domain [8, 5, 19, 20, 53, 61, 63], thus requiring scaling of their depth prediction. However, due to the scale-ambiguity of multi-view photometric objective, such models are normally evaluated by aligning predictions to ground-truth at test time (typically median-scaling  [16, 17, 85]), at the expense of practicality since ground-truth might not be feasible during real-world application. To enable generalization across different scenes, some (semi-) supervised depth models trained with single images adopt image-level normalization techniques to generate affine-invariant depth representations (i.e. relative depth) [23, 42, 43, 66, 72]. HND [72] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [66] learns from large-scale automatically annotated data. DPT [42] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [43] mixes multiple datasets with training objectives invariant to depth range and scale. Marigold [23] associates fine-tuning protocol with a diffusion model. However, by definition, relative depth is scaleless, which limits their use in applications that require metric scaled reconstruction. Our proposed RSAaddresses this limitation by grounding relative depth into a metric scale, enabling metric scale 3D reconstruction.
Relative to metric depth transfer.
To transform the predicted relative depth into metric depth for evaluation and real-world application, ZoeDepth [3] fine-tunes a metric bins module on each metric depth dataset to output metric depth. Depth Anything [66] follows ZoeDepth [3] using a decoder where a metric bins module computes per-pixel depth bin centers that are linearly combined to output the metric depth. MiDas [43] and Marigold [23] use linear fit to align predictions and ground truth in scale and shift for each image in inverse-depth space based on the least-squares criterion before measuring errors. DPT [42] fine-tunes a global scale and shift on metric depth datasets. DistDepth [61] conducts transfer by leveraging left-right stereo consistency to integrate metric scale into a scale-agnostic depth network. ZeroDepth [20] achieves transformation using input-level geometric embedding to learn an indoor scale prior over objects via a variational latent representation. However, metric decoders are limited to specific datasets, and aligning scale and shift of predictions requires ground truth during test time. RSA produces scale using text to transfer relative depth to metric depth across domains and does not require ground truth during test time.
@@ 3 Method
RSA. To infer the parameters of the linear transformation to align relative depth to metric scale, we employ the existing pretrained CLIP text encoder [41] as a feature extractor. Having been trained on a large scale dataset, CLIP offers an latent space suitable to preprocess object-centric text descriptions. We note that CLIP text encoder is frozen within RSA. Given text descriptions $\textbf{t}=\{t_{1},t_{2},...\}$, we first encode them into text embeddings and feed them into a 3-layer shared multi-layer perceptron (MLP) to project them into $k=256$ hidden dimensions followed by two separate sets of 3-layer MLPs, one serves as the output head $\psi_{\hat{\alpha}}:\mathbb{R}^{k}\mapsto\mathbb{R}$ for scale parameter $\hat{\alpha}$ and the other as the output head $\psi_{\hat{\beta}}:\mathbb{R}^{k}\mapsto\mathbb{R}$ for shift $\hat{\beta}$ parameter. For ease of notation, we refer to $\psi$ as the parameters for the shared MLP as well as the output heads, where $(\hat{\alpha},\hat{\beta})=g_{\psi}(\textbf{t})$.
@@
Text prompt design. To test our hypothesis, we require text descriptions to be paired with images. As standard benchmarks do not provide the text description of each image, we extend existing datasets by associating each image with several text descriptions. To achieve this, we propose to use off-the-shelf models to generate different kinds of text. First, we considered structured text, which adheres to a certain template. We use a panoptic segmentation model MaskDINO [30] to extract the significant objects and background in the image. For an input image $I$, the segmentation model returns a set of $B$ object and background instance $\{\textbf{n}^{(1)},\textbf{c}^{(i)}\}_{i=1}^{B}$, where $\textbf{c}^{(i)}$ denotes the class of the object or background, and $\textbf{n}^{(i)}$ denotes the number of instances of the object or background. Using the set of instances, a structured caption for an image $I$ can be obtained: “ An image with $\textbf{n}^{(1)}$ $\textbf{c}^{(1)}$, $\textbf{n}^{(2)}$ $\textbf{c}^{(2)}$, …, $\textbf{n}^{(B)}$ $\textbf{c}^{(B)}$. ” We will shuffle the order of instances to produce 5 different structured captions for each image. Then we consider the natural text, where the text doesn’t adhere to certain templates and is closer to human descriptions. We use a visual question answering model LLaVA v1.6 Vicuna and LLaVA v1.6 Mistral [35]. For each model, we prompt it with the input image and a prompt, asking the model to describe the image. For each model, we provide 5 different prompts to produce different natural captions. During training, in each iteration, for a given image, we randomly select one caption from those 15 captions to predict scale and shift.
@@
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\log_{10}\downarrow$ RMSE $\downarrow$ ZoeDepth Image NYUv2 0.951 0.994 0.999 0.077 0.033 0.282 DistDepth DA NYUv2 0.706 0.934 - 0.289 - 1.077 DistDepth DA,Median NYUv2 0.791 0.942 0.985 0.158 - 0.548 ZeroDepth DA - 0.901 0.961 - 0.100 - 0.380 ZeroDepth DA,Median - 0.926 0.986 - 0.081 - 0.338 DPT Median NYUv2 0.736 0.919 0.981 0.181 0.073 0.912 Linear Fit NYUv2 0.926 0.991 0.999 0.094 0.040 0.332 Global NYUv2 0.904 0.988 0.998 0.109 0.045 0.357 Image NYUv2 0.914 0.990 0.998 0.097 0.042 0.350 Image NYUv2,KITTI 0.911 0.989 0.998 0.098 0.043 0.355 RSA (Ours) NYUv2 0.916 0.990 0.998 0.097 0.042 0.347 RSA (Ours) NYUv2,KITTI 0.913 0.988 0.998 0.099 0.042 0.352 MiDas Median NYUv2 0.449 0.694 0.850 0.411 0.151 2.010 Linear Fit NYUv2 0.780 0.970 0.995 0.151 0.069 0.433 Global NYUv2 0.689 0.949 0.992 0.183 0.078 0.600 Image NYUv2 0.729 0.958 0.994 0.175 0.072 0.563 Image NYUv2,KITTI 0.724 0.952 0.992 0.173 0.074 0.579 RSA (Ours) NYUv2 0.731 0.955 0.993 0.171 0.072 0.569 RSA  (Ours) NYUv2,KITTI 0.737 0.959 0.993 0.168 0.071 0.561 DepthAnything Median NYUv2 0.480 0.734 0.886 0.353 0.135 1.743 Linear Fit NYUv2 0.965 0.993 0.997 0.058 0.025 0.232 Global NYUv2 0.630 0.926 0.987 0.199 0.087 0.646 Image NYUv2 0.749 0.965 0.997 0.169 0.068 0.517 Image NYUv2,KITTI 0.710 0.947 0.992 0.181 0.075 0.574 RSA (Ours) NYUv2 0.775 0.975 0.997 0.147 0.065 0.484 RSA (Ours) NYUv2,KITTI 0.776 0.974 0.996 0.148 0.065 0.498
@@
Table 1: Quantitative results on NYUv2. Global refers to a single scale and shift for entire dataset. Image denotes predicting scale and shift from an image. Red denotes scaling that uses ground truth. Median scaling per image is a common practice for evaluation. Linear fit denotes fitting to ground truth for each image. DA refers to domain adaptation. ZoeDepth performs per-pixel refinement. RSA (yellow) when trained with NYUv2 and KITTI generalizes better than using image to predict the transformation parameters.
@@ 4 Experiments
Datasets. We present our main result on two datasets: NYUv2  [45] for indoor scenes, and KITTI [15] for outdoor scenes. NYUv2 contains 480$\times$640 images with depth values from $1\times 10^{-3}$ to 10 meters. We follow [28, 34, 78] for the dataset partition, which contains 24,231 train images and 654 test images. KITTI contains 352$\times$1216 images where depth values from $1\times 10^{-3}$ to 80 meters. We adopt the Eigen Split [9] consisting of 23,488 training images and 697 testing images. Following [2, 70], we remove samples without valid ground truth, leaving 652 valid images for testing. We also report zero-shot generalization results on SUN-RGBD [47], which contains 5050 testing images.
Depth models. For DPT [42], we use DPT-Hybrid fine-tuned for NYUv2 and KITTI respectively, with 123M parameters. For MiDas [43], we use MiDaS 3.1 Swin2_large-384 with 213M parameters. For DepthAnything [66] (DA), we use Depth-Anything-Small with 24.8M parameters. Different from our setting, DepthAnything evaluates using ZoeDepth [3]’s depth decoder that is separately fine-tuned on NYUv2 and KITTI to produce a pixel-wise scale, and MiDas evaluates by aligning prediction with ground truth in scales and shifts. We re-implement several baselines aligning with the setting of predicting a global scale and shift for scaling relative depth, provided in Table 1 and  2.
Hyperparameters. We use the Adam [24] optimizer without weight decay. The learning rate is reduced from $3\times 10^{-5}$ to $1\times 10^{-5}$ by a cosine learning rate scheduler. The model is trained for 200 epochs under this scheduler. We run our experiment on GeForce RTX 3090 GPUs, with 24GB memory. For reference, if using a single GPU, the training time for DepthAny-RSA on jointly NYUv2 and KITTI takes roughly 1 week (6.46 days).
@@
Quantitative results. We show results on NYUv2 in Table 1, and KITTI in Table 2, where we improve over baselines across all evaluation metrics and approach the performance of using ground truth for scaling. Following DPT, we optimize the scale and shift for global scaling baseline over the training set and use it for evaluation (i.e. same scale and shift for all test data). We obtain the “Image" baseline by substituting CLIP text features with CLIP image features. Following  [43], we perform a linear regression to find the scale and shift that minimizes the least-square error between the ground truth metric depth and the predicted metric depth. Additionally, we also test median scaling [17], a common practice for evaluation. Both linear fitting and median scaling are shown to demonstrate what is achievable if one were to directly fit to ground truth. We train for separated RSA models for each dataset, as well as a unified RSA model combining both KITTI and NYUv2.
Although RSA trained on a single dataset may achieve slightly better performance than the unified model, the difference is minimal. In some metrics, the unified model even outperforms, demonstrating the generalizability of using language as input, given the narrow domain gap for language. Additionally, RSA models consistently outperform “Image" baselines in both in-domain and cross-domain scenarios and are comparable with existing methods under various settings. Our method significantly narrows the performance gap to that of using ground truth for scaling, validating the effectiveness of using language instead of the input image to predict scale. Figure 5 shows a curve fitting plotted for our predicted (inverse) scale against median scale. The trend line shows that the scale inferred from language matches well to the median of the ground truth, which is a robust estimator.

### v2

@@ Abstract
We propose a method for metric-scale monocular depth estimation. Inferring depth from a single image is an ill-posed problem due to the loss of scale from perspective projection during the image formation process. Any scale chosen is a bias, typically stemming from training on a dataset; hence, existing works have instead opted to use relative (normalized, inverse) depth. Our goal is to recover metric-scaled depth maps through a linear transformation. The crux of our method lies in the observation that certain objects (e.g., cars, trees, street signs) are typically found or associated with certain types of scenes (e.g., outdoor). We explore whether language descriptions can be used to transform relative depth predictions to those in metric scale. Our method, RSA, takes as input a text caption describing objects present in an image and outputs the parameters of a linear transformation which can be applied globally to a relative depth map to yield metric-scaled depth predictions. We demonstrate our method on recent general-purpose monocular depth models on indoors (NYUv2, VOID) and outdoors (KITTI). When trained on multiple datasets, RSA can serve as a general alignment module in zero-shot settings. Our method improves over common practices in aligning relative to metric depth and results in predictions that are comparable to an upper bound of fitting relative depth to ground truth via a linear transformation. Code is available at: https://github.com/Adonis-galaxy/RSA
@@ 1 Introduction
We consider whether an additional modality can be used to resolve the scale ambiguity in single-image 3D reconstruction, i.e., transforming scaleless relative depth to metric depth. One might observe that natural (including man-made) scenes do not occur by chance, but rather by design, with the regularity of object-scene co-occurrences [27, 38]: Certain scenes (e.g., outdoors) are composed of certain categories of objects (e.g., cars, trees, buildings) and associated with a certain order of magnitude in scale (e.g. tens of meters). Hence, we hypothesize that language, in the form of text captions or descriptions, can be used to infer the scale of the 3D scene and to transform relative depth to metric depth. The choice of language also has practical value in that it does not require costly data acquisition with an additional synchronized and calibrated sensor (e.g., lidar, time-of-flight). With publicly available pre-trained panoptic segmentation or object detection models, image captioners, and vision-language models, one can automate the data acquisition, training, and inference process. Nonetheless, they are not a necessary part of the work, but may facilitate ease of use, to simulate the language description provided by human users in practical scenarios.
@@
To test the feasibility of our hypothesis, we consider monocular depth estimation, where a strong prior is necessary for inference; this prior may come from an image, or an independent modality such as language. Specifically, we consider monocular depth models belonging to the general-purpose, relative depth estimation paradigm to control for side effects such as the scale learned along with shapes in existing works that focus on predicting metric depth for a specific dataset. To this end, we assume access to a pre-trained scaleless monocular depth estimation model; aside from a set of images, we also assume text captions describing them. As the standard procedure to factor out scale is to normalize through a linear transformation, we propose to learn a parameterized function that predicts the parameters of a linear transformation based on the text description. Applying them to a relative depth map and inverting its values will transform it to a metric-scaled depth map to resolve the scale ambiguity. We term our method RSA, as an acronym for “Resolving Scale Ambiguities”.
@@
In one mode, like existing works that also transform relative to metric depth, albeit with images and domain-specific scaling [4], we train specific models of RSA for specific datasets (e.g., NYUv2 [59], KITTI [22], VOID [73]). In another mode, when trained on multiple datasets across different domains, RSA generalizes well and can not only handle images sampled from the datasets it was trained on, but also those from novel datasets in a zero-shot manner. The use of language, which is invariant to illumination, object orientation, occlusion, scene layout, etc., many of the nuisance variables that vision algorithms are sensitive to, demonstrates a promising avenue for general-purpose relative to metric scale recovery to complement the growing works in monocular depth estimation. We evaluate our method on indoor and outdoor benchmarks, where we improve over common practices in aligning relative depth to metric scale. We show that using RSA is comparable to matching relative depth via a linear transformation to the ground truth, or scaling using the median value of the ground truth, both of which are considered oracle relative to metric recovery methods.
@@ 2 Related Work
Monocular depth estimation using relative depth. Trained across different domains, self-supervised depth estimators trained with multi-view photometric objects produce up-to-scale predictions that are linearly correlated with their absolute depth values across the domain [10, 6, 30, 31, 69, 78, 81], thus requiring scaling of their depth prediction. However, due to the scale-ambiguity of multi-view photometric objective, such models are normally evaluated by aligning predictions to ground-truth at test time (typically median-scaling  [24, 26, 103]), at the expense of practicality since ground-truth might not be feasible during real-world application. To enable generalization across different scenes, some (semi-) supervised depth models trained with single images adopt image-level normalization techniques to generate affine-invariant depth representations (i.e. relative depth) [34, 56, 57, 84, 90]. HND [90] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [84] learns from large-scale automatically annotated data. DPT [56] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [57] mixes multiple datasets with training objectives invariant to depth range and scale. Marigold [34] associates fine-tuning protocol with a diffusion model. However, by definition, relative depth is scaleless, which limits their applications that require metric scaled reconstruction. Our proposed RSA addresses this limitation by grounding relative depth into a metric scale, enabling metric scale 3D reconstruction.
Relative to metric depth transfer. To transform the predicted relative depth into metric depth for evaluation and real-world application, ZoeDepth [4] fine-tunes a metric bins module on each metric depth dataset to output metric depth. Depth Anything [84] follows ZoeDepth [4] using a decoder where a metric bins module computes per-pixel depth bin centers that are linearly combined to output the metric depth. MiDas [57] and Marigold [34] use linear fit to align predictions and ground truth in scale and shift for each image in inverse-depth space based on the least-square criterion before measuring errors. DPT [56] fine-tunes a global scale and shift on metric depth datasets. DistDepth [78] conducts transfer by leveraging left-right stereo consistency to integrate metric scale into a scale-agnostic depth network. ZeroDepth [31] achieves transformation using input-level geometric embedding to learn an indoor scale prior over objects via a variational latent representation. However, metric decoders are limited to specific datasets, and aligning scale and shift of predictions requires ground truth during test time. RSA produces scale using text to transfer relative depth to metric depth across domains and does not require ground truth during test time.
@@ 3 Method
RSA. To infer the parameters of the linear transformation to align relative depth to metric scale, we employ the existing pretrained CLIP text encoder [55] as a feature extractor. Having been trained on a large scale dataset, CLIP offers an latent space suitable to preprocess object-centric text descriptions. We note that CLIP text encoder is frozen within RSA. Given text descriptions $\textbf{t}=\{t_{1},t_{2},...\}$, we first encode them into text embeddings and feed them into a 5-layer shared multi-layer perceptron (MLP) to project them into $k=256$ hidden dimensions followed by two separate sets of 5-layer MLPs, one serves as the output head $\psi_{\hat{\alpha}}:\mathbb{R}^{k}\mapsto\mathbb{R}_{+}$ for scale parameter $\hat{\alpha}$ and the other as the output head $\psi_{\hat{\beta}}:\mathbb{R}^{k}\mapsto\mathbb{R}_{+}$ for shift $\hat{\beta}$ parameter. Scale and shift are assumed to be positive in favor of optimization. For ease of notation, we refer to $\psi$ as the parameters for the shared MLP as well as the output heads, where $(\hat{\alpha},\hat{\beta})=g_{\psi}(\textbf{t})$.
@@
Text prompt design. To test our hypothesis, we require text descriptions to be paired with images. As standard benchmarks do not provide the text description of each image, we extend existing datasets by associating each image with several text descriptions. To achieve this, we propose to use off-the-shelf models to generate different kinds of text. First, we considered structured text, which adheres to a certain template. We use a panoptic segmentation model MaskDINO [41] to extract the significant objects and background in the image. For an input image $I$, the segmentation model returns a set of $B$ object and background instances $\{\textbf{n}^{(i)},\textbf{c}^{(i)}\}_{i=1}^{B}$, where $\textbf{c}^{(i)}$ denotes the class of the object or background, and $\textbf{n}^{(i)}$ denotes the number of instances of the object or background. Using the set of instances, a structured caption for an image $I$ can be obtained: “ An image with $\textbf{n}^{(1)}$ $\textbf{c}^{(1)}$, $\textbf{n}^{(2)}$ $\textbf{c}^{(2)}$, …, $\textbf{n}^{(B)}$ $\textbf{c}^{(B)}$. ” We will shuffle the order of instances to produce 5 different structured captions for each image. Then we consider the natural text, where the text doesn’t adhere to certain templates and is closer to human descriptions, we use two visual question-answering models LLaVA v1.6 Vicuna and LLaVA v1.6 Mistral [47]. For each model, we prompt it with the input image and a prompt, asking the model to describe the image. For each model, we provide 5 different prompts to produce different natural captions. During training, in each iteration, for a given image, we randomly select one caption from those 15 captions to predict scale and shift.
@@
Models Scaling Dataset $\delta<1.25\uparrow$ $\delta<1.25^{2}\uparrow$ $\delta<1.25^{3}\uparrow$ Abs Rel $\downarrow$ $\log_{10}\downarrow$ RMSE $\downarrow$ ZoeDepth Image NYUv2 0.951 0.994 0.999 0.077 0.033 0.282 DistDepth DA NYUv2 0.706 0.934 - 0.289 - 1.077 DistDepth DA,Median NYUv2 0.791 0.942 0.985 0.158 - 0.548 ZeroDepth DA - 0.901 0.961 - 0.100 - 0.380 ZeroDepth DA,Median - 0.926 0.986 - 0.081 - 0.338 DPT Median NYUv2 0.736 0.919 0.981 0.181 0.073 0.912 Linear Fit NYUv2 0.926 0.991 0.999 0.094 0.040 0.332 Global NYUv2 0.904 0.988 0.998 0.109 0.045 0.357 Image NYUv2 0.914 0.990 0.998 0.097 0.042 0.350 Image NYUv2,KITTI 0.911 0.989 0.998 0.098 0.043 0.355 Image NYUv2,KITTI,VOID 0.903 0.985 0.997 0.100 0.045 0.367 RSA (Ours) NYUv2 0.916 0.990 0.998 0.097 0.042 0.347 RSA (Ours) NYUv2,KITTI 0.913 0.988 0.998 0.099 0.042 0.352 RSA (Ours) NYUv2,KITTI,VOID 0.912 0.989 0.998 0.099 0.043 0.355 MiDas Median NYUv2 0.449 0.694 0.850 0.411 0.151 2.010 Linear Fit NYUv2 0.780 0.970 0.995 0.151 0.069 0.433 Global NYUv2 0.689 0.949 0.992 0.183 0.078 0.600 Image NYUv2 0.729 0.958 0.994 0.175 0.072 0.563 Image NYUv2,KITTI 0.724 0.952 0.992 0.173 0.074 0.579 Image NYUv2,KITTI,VOID 0.712 0.948 0.988 0.181 0.075 0.583 RSA (Ours) NYUv2 0.731 0.955 0.993 0.171 0.072 0.569 RSA  (Ours) NYUv2,KITTI 0.737 0.959 0.993 0.168 0.071 0.561 RSA (Ours) NYUv2,KITTI,VOID 0.709 0.944 0.989 0.173 0.076 0.580 DepthAnything Median NYUv2 0.480 0.734 0.886 0.353 0.135 1.743 Linear Fit NYUv2 0.965 0.993 0.997 0.058 0.025 0.232 Global NYUv2 0.630 0.926 0.987 0.199 0.087 0.646 Image NYUv2 0.749 0.965 0.997 0.169 0.068 0.517 Image NYUv2,KITTI 0.710 0.947 0.992 0.181 0.075 0.574 Image NYUv2,KITTI,VOID 0.702 0.943 0.990 0.178 0.078 0.583 RSA (Ours) NYUv2 0.775 0.975 0.997 0.147 0.065 0.484 RSA (Ours) NYUv2,KITTI 0.776 0.974 0.996 0.148 0.065 0.498 RSA (Ours) NYUv2,KITTI,VOID 0.752 0.964 0.992 0.156 0.071 0.528
@@
Table 1: Quantitative results on NYUv2. RSA (yellow), especially when trained with multiple datasets, generalizes better than using images to predict the transformation parameters. Global refers to optimizing a single scale and shift for the entire dataset (same scale and shift for every sample). Image denotes predicting scales and shifts using images. Red denotes scaling that uses ground truth. Median indicates scaling using the ratio between median of depth prediction and ground truth. Linear fit denotes optimizing scale and shift to fit to ground truth for each image. DA refers to domain adaptation. ZoeDepth performs per-pixel refinement.
@@ 4 Experiments
Datasets. We present our main result on three datasets: NYUv2  [59] and VOID [73] for indoor scenes, and KITTI [22] for outdoor scenes. NYUv2 contains images with a resolution of 480$\times$640 where depth values from $1\times 10^{-3}$ to 10 meters. We follow [39, 46, 96] for the dataset partition, which contains 24,231 train images and 654 test images. VOID contains images with a resolution of 480$\times$640 where depth values from 0.2 to 5 meters. It contains 48,248 train images and 800 test images following the official splits [73]. KITTI contains images with a resolution of 352$\times$1216 where depth values from $1\times 10^{-3}$ to 80 meters. We adopt the Eigen Split [13] consisting of 23,488 training images and 697 testing images. Following [3, 88], we remove samples without valid ground truth, leaving 652 valid images for testing. We also report zero-shot generalization results on SUN-RGBD [62], which contains 5050 testing images, and DDAD [28], which contains 3950 validation images.
Depth models. For DPT [56], we use DPT-Hybrid fine-tuned for NYUv2 and KITTI respectively, with 123M parameters. We used the one fine-tuned on NYUv2 for VOID. For MiDas [57], we use MiDaS 3.1 Swin2_large-384 with 213M parameters. For DepthAnything [84], we use Depth-Anything-Small with 24.8M parameters. Different from our setting, DepthAnything evaluates using ZoeDepth [4]’s depth decoder that is separately fine-tuned on NYUv2 and KITTI to produce a pixel-wise scale, and MiDas evaluates by aligning prediction with ground truth in scales and shifts. We re-implement several baselines aligning with the setting of predicting a global scale and shift for scaling relative depth, provided in Table 1 and  2.
Hyperparameters. We use the Adam [35] optimizer without weight decay. The learning rate is reduced from $3\times 10^{-5}$ to $1\times 10^{-5}$ by a cosine learning rate scheduler. The model is trained for 50 epochs under this scheduler. We run our experiment on GeForce RTX 3090 GPUs, with 24GB memory. For reference, if using a single GPU, the training time for RSA with DepthAnything on jointly NYUv2, KITTI, and VOID for 50 epochs takes 57 hours..